Find the maximum and minimum values of f(x, y, z) = xyz subject to the constraint 4x2+9y+z2 = 4 Assume that y ≥ 0. Why is this assumption required?


1
Expert's answer
2021-11-09T16:28:40-0500

"\displaystyle
\text{First we define a new a function F(x,y)}\\
F(x,y,z,\lambda) = xyz + \lambda(4x^2+9y+z^2)\\
\text{Next, we differentiate F with respect to x,y and z and equate them}\\
\text{to 0}\\
F_x(x,y,z,\lambda) = yz + 8x\lambda=0 \qquad -(1)\\
F_y(x,y,z,\lambda) = xz + 9\lambda=0 \qquad-(2)\\
F_z(x,y,z,\lambda) = xy+ 2z\lambda=0 \qquad -(3)\\
\text{Using elimination method, we obtain }\\
z=2x, y = \frac{8}{9}x^2\\
\text{Substituting in the given constraint, we have that}\\
x=\pm \frac{1}{2}, y= \pm \frac{2}{9}, \pm 1\\
f(\frac{1}{2},\frac{2}{9},1)=\frac{1}{9}, f(-\frac{1}{2},-\frac{2}{9},-1)=-\frac{1}{9}\\\text{Therefore, $\frac{1}{9}$ is an absolute maximum at point ($\frac{1}{2},\frac{2}{9},1)$ and}\\
\text{point $-\frac{1}{9}$ is an absolute maximum at point ($-\frac{1}{2},-\frac{2}{9},-1)$}"